OpenStudy (brittany):

Write the equation in slope-intercept form of the line that has a slope of 6 and contains the point (1, 1).

OpenStudy (anonymous):

y = 6x + 1

OpenStudy (brittany):

incorrect

OpenStudy (anonymous):

yea i know...hang on

OpenStudy (brittany):

ok thanks :)

OpenStudy (anonymous):

y - 1 = 6( x -1) y - 1 = 6x - 6 y = 6x -5 slope is 6 y intercept is -5

OpenStudy (anonymous):

equation in slope intercept form is y = 6x -5 and that is correct

OpenStudy (brittany):

thanks :)

OpenStudy (anonymous):

no problem